OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4? A. y = –3x + 4 B. y = –3x – 4 C. y = 3x + 4 D. y = 3x – 4

OpenStudy (anonymous):

B

OpenStudy (the_fizicx99):

y = mx + b m = slope b = y intercept y = -3x - 4

OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 3x + 7y = 11? A. (6, –1) B. (1, –2) C. (0,–4) D. (0, 4)

OpenStudy (the_fizicx99):

Plug int those values for x and y

OpenStudy (the_fizicx99):

3(0) + 7(4) = 11 28 = 11 Its not D :p